OpenStudy (anonymous):

If ‘x’ number of children are playing in snow and each one throws a snowball to other then, no. of snow balls thrown are

OpenStudy (stamp):

\[\frac{snowballs}{child}children=thrown\ snowballs\]

hartnn (hartnn):

if 2 children are playing then no. of snow balls thrown = 1+1 =2 = 2*(1) if 3 children are playing then no. of snow balls thrown = 3(2) = 6 [A,B,C then A will throw at B and C ---> 2 for A ....2 for B, 2 for C--->total=6] if 4 children are playing then no. of snow balls thrown = 4*(3) = 12 so, in general, if x children are playing then no. of snow balls thrown =x (x-1)